# Towards Cross-modal Backward-compatible Representation Learning for Vision-Language Models

### v1

@@ 1 Introduction
With these generated synthetic old image and text embeddings, which can be considered as aligned with the new embeddings, we aim to fine-tune the new model. In this process, far fewer supervised image-text pairs are required than in the original training dataset of the VLP. This also allows us to avoid using old model parameters during training, thereby enhancing training efficiency. In addition, we incorporate parameter-efficient fine-tuning schemes [19, 12] into the new model, which add small trainable parameters and do not harm any of the new model’s original parameters during training. This not only accelerates the new model’s fine-tuning process but also allows for an easy reversion to the original new-to-new retrieval by simply removing the additional parameters.
@@ 2 Related Works
Backward-compatible Training. The concept of Backward-compatible Training (BT) was first introduced in the study [33]. This approach influences a new model with the learned classifier of the old model, thereby achieving backward compatibility between the old and new models. However, BT can degrade the original representational performance of the new model. To overcome this, [25] proposed aligning class-wise centers presented by the old and new models. Another approach to achieve backward compatibility is to use an additional projection to map the old embedding into the new embedding space by adding a lightweight transformation, as suggested in [13, 37]. The work of [28] further adds an auxiliary feature in preparation for future updates, while [46] uses additional dimensionality in the embedding to obtain compatibility. An online strategy that backfills the gallery on the fly is explored in [42], and [41] addresses the model regression problem. Despite this progress, the scenario considering cross-modal retrieval between image and text, which has many real-world applications, remains unexplored. Our study on XBT in this paper aims to fill this gap.
@@
Vision-Language Continual and Transfer Learning The fields of continual learning [2, 5, 7] and transfer learning [24, 47] bear similarities to backward-compatible representation learning, as all aim to update an existing model to boost performance. In the realm of multi-modal representation learning for cross-modal retrieval, continual learning approaches like [39] propose methods to prevent catastrophic forgetting across different modalities. Transfer learning approaches like [44] suggest strategies for transferring knowledge from previously labeled categories (source modality) to new, unlabeled categories (target modality). However, our proposed XBT stands apart from these methods as it specifically tackles the challenge of maintaining backward compatibility between old and new models. This unique attribute makes XBT ideally suited for backfill-free retrieval scenarios, where the embeddings of the old model remain unchanged, yet we can still leverage the enhanced performance of the new model.
3 Methodology
@@
Our goal is to construct a backfill-free, embedding-based, cross-modal retrieval system using VLP models, which are configured with an image encoder $E^{I}$ and a text encoder $E^{T}$. When a new better performing VLP model $\{E^{I}_{new},E^{T}_{new}\}$ emerges, we aim to ensure its compatibility with the old model $\{E^{I}_{old},E^{T}_{old}\}$ that was used to construct the gallery. To achieve this, we introduce Cross(X)-modal Backward-compatible Training (XBT). Given an image $x$ and text caption $t$, XBT enables retrieval between a new image embedding $v_{new}=E^{I}_{new}(x)$ and the text embeddings ($w_{old}$), as well as between a new text embedding $w_{new}=E^{T}_{new}(t)$ and the image embeddings ($v_{old}$). We denote backward compatible embeddings as $\bar{v}$ and $\bar{w}$. All embeddings we utilize in this work are l2-normalized.
@@ 3.1 Criterion for Cross-modal Backward Compatibility
$d(w_{new_{i}},v_{old_{j}})\leq d(w_{old_{i}},v_{old_{j}}),\forall y_{i}=y_{j},d(w_{new_{i}},v_{old_{j}})\geq d(w_{old_{i}},v_{old_{j}}),\forall y_{i}\neq y_{j},$ (1)
$d(v_{new_{i}},w_{old_{j}})\leq d(v_{old_{i}},w_{old_{j}}),\forall y_{i}=y_{j},d(v_{new_{i}},w_{old_{j}})\geq d(v_{old_{i}},w_{old_{j}}),\forall y_{i}\neq y_{j},$
@@ 3.2 Text-only Pretraining
Figure 2: An illustration of the text-only pretraining of $\phi$ (above), and XBT with $\phi$ (below). Using only text samples, $\phi$ is trained to approximate distribution of old text embeddings from that of new ones. During training, noise $\epsilon$ is injected to input of $\phi$. After pretraining, the same $\phi$ is used to generate both of synthetic old image and text embeddings from the new VLP embeddings and train to learn cross-modal backward-compatible representation.
@@ 4.1 Setup
Model Training. Our XBT process involves two distinct training stages: text-only pretraining (See Sec. 3.2) and image-text supervised training (See Sec. 3.3). For the text-only pretraining stage, we utilize the text samples from a subset of the 115M filtered web-collected image-text paired dataset from BLIP [20], comprising around 67M available pairs (58.2% of the total). We construct a text corpus $D_{T}$ in Eqn. 3 using these text samples. For the subsequent image-text supervised training stage, we use a smaller subset of 4M image-text pairs from the same dataset to construct a supervised dataset $D$ in Eqn. 5. It is important to note that these subsets are significantly smaller than the dataset scale (400M, 2B or more) used to build CLIP models [27, 31].
Model Evaluation. We validate the effectiveness of XBT with a benchmark comprising three popular image-text paired datasets for cross-modal retrieval evaluation. The first is the nocaps dataset [1], which offers a diverse distribution of object names and facilitates a more detailed analysis. We employ the validation split of this dataset, which consists of images each paired with 10 relevant textual captions, totaling 4,500 images and 45,000 captions.
To evaluate XBT at a larger scale, we also include the Flickr [14] and COCO [22] datasets, which consist of images that are each paired with 5 relevant textual captions. For the Flickr dataset, we use the entire dataset, encompassing 31,783 images and 158,915 captions. For the COCO dataset, we use the validation split, which includes 35,136 images and 175,680 captions. We notate the two datasets as Flickr-31K and COCO-35K respectively for the remainder of the paper.
@@ 4.2 Main Results
Table 2: Cross-modal retrieval results on Flickr-31K.
Old Model / New Model Method Text $\text{Text Query}(w,\bar{w})/{\text{Image Gallery}}(v)$ Image $\text{Image Query}(v,\bar{v})/{\text{Text Gallery}}(w)$ Case R@1 R@5 R@10 R@50 Case R@1 R@5 R@10 R@50 Original CLIP-ViT-B32 - $w/v$ 21.54 41.24 50.72 72.69 $v/w$ 40.35 64.42 73.39 89.91 Cross-modal Backward Compatible Training CLIP-ViT-B32 / CLIP-ViT-L14 Base 19.02 37.25 46.63 69.32 36.74 60.61 70.22 88.34 XBT $\bar{w}_{new}/v_{old}$ 22.38 42.71 52.41 74.80 $\bar{v}_{new}/w_{old}$ 42.47 66.04 74.87 90.79 Base 32.29 55.18 64.68 83.60 36.10 59.80 69.95 89.01 XBT $\bar{w}_{new}/\bar{v}_{new}$ 39.59 62.89 71.75 88.13 $\bar{v}_{new}/\bar{w}_{new}$ 43.50 68.50 77.99 93.21 CLIP-ViT-B32 / CLIP-ViT-H14 Base 20.27 39.17 48.59 70.67 35.53 59.44 69.06 87.67 XBT $\bar{w}_{new}/v_{old}$ 23.70 44.47 54.19 76.03 $\bar{v}_{new}/w_{old}$ 41.09 65.18 74.32 90.16 Base 32.94 55.94 65.26 83.77 34.81 59.55 69.96 88.81 XBT $\bar{w}_{new}/\bar{v}_{new}$ 40.42 63.90 72.74 88.47 $\bar{v}_{new}/\bar{w}_{new}$ 46.35 71.18 80.01 94.08
Table 3: Cross-modal retrieval results on COCO-35K.
Old Model / New Model Method Text $\text{Text Query}(w,\bar{w})/{\text{Image Gallery}}(v)$ Image $\text{Image Query}(v,\bar{v})/{\text{Text Gallery}}(w)$ Case R@1 R@5 R@10 R@50 Case R@1 R@5 R@10 R@50 Original CLIP-ViT-B32 - $w/v$ 14.44 30.20 38.96 62.30 $v/w$ 28.62 50.17 59.67 80.22 Cross-modal Backward Compatible Training CLIP-ViT-B32 / CLIP-ViT-L14 Base 12.97 28.14 36.48 59.78 26.32 46.91 56.54 77.86 XBT $\bar{w}_{new}/v_{old}$ 15.55 32.27 41.34 65.11 $\bar{v}_{new}/w_{old}$ 30.73 52.58 62.30 81.83 Base 22.16 41.65 50.98 73.06 27.26 48.16 57.80 79.57 XBT $\bar{w}_{new}/\bar{v}_{new}$ 27.61 48.69 58.08 78.83 $\bar{v}_{new}/\bar{w}_{new}$ 32.92 55.37 64.89 84.67 CLIP-ViT-B32 / CLIP-ViT-H14 Base 13.88 29.55 38.23 61.57 26.19 46.75 56.30 77.96 XBT $\bar{w}_{new}/v_{old}$ 15.82 32.79 41.84 65.45 $\bar{v}_{new}/w_{old}$ 30.33 51.28 59.97 81.66 Base 23.20 43.22 52.70 74.68 27.31 48.56 58.60 80.09 XBT $\bar{w}_{new}/\bar{v}_{new}$ 28.14 49.48 58.88 79.32 $\bar{v}_{new}/\bar{w}_{new}$ 34.54 57.29 66.79 85.92
Large Scale Retrieval. In this experiment, we assess the performance of cross-modal backward compatibility in more practical use cases. We use larger datasets, namely Flickr-31K and COCO-35K, as shown in Tables 2 and 3. These datasets provide more than six times the number of samples typically used in image-text retrieval literature [36, 6], which make the retrieval process more challenging. We use the second-best performing Base method from Table 1 as a basis for comparison. To validate the applicability of XBT across various VLP models, we utilize three CLIP baselines: B32, L14, and H14. Additionally, we explore different combinations, such as B32/L14 and B32/H14, to further analyze the effectiveness of XBT, beyond the comparisons made in Table 1 (B32/L14 and L14/H32).
@@ 4.3 Further Analysis
[…]
@@
Ablation Study. To validate XBT, we conduct further analysis as shown in Table 5. By comparing (a) and (b), we observe that introducing noise during $\phi$ training aids in generalization. The comparison between (c) and (d) demonstrates that the scale of $D_{T}$ is important, supporting our assumption that a sufficient number of text samples can help build a robust $\phi$. (e) outperforms (a), confirming that utilizing more image-text pairs can enhance XBT. The lower performance of (f) also aligns with the notion that the number of image-text pairs is crucial. In (g), when we replace $D$ with CC3M [32], from which we can obtain around 2.4M, the performance is similar to (f), suggesting that XBT can be effectively applied with other datasets.
Different VLP models. To further explore the capacity of the VLP model architecture’s generalization of XBT, we evaluate it using BLIP [20] Base and Large models. We employ checkpoints of Salesforce/blip-itm-base-coco and Salesforce/blip-itm-large-coco from the Huggingface library. We apply XBT on old and new BLIP models in the same fashion with our CLIP applications, and the results appear in Table 5. From the results, we confirm that XBT provides cross-modal backward-compatibility to the BLIP models too.
5 Discussion & Conclusion
Potential Broader Impact and Limitation. This paper highlights our efforts to enhance the field of Machine Learning, particularly in the area of multi-modal embedding-based representation learning. Although our work may have societal implications, we do not believe there are any that require specific emphasis in this context. A potential limitation of the XBT system is that, despite the efficient learning approach reducing the need for image-caption pairs, its performance may still be limited by the quality, diversity, and representational richness of the data during training.
Conclusion. In this paper, we introduced Cross-modal Backward-compatible Training (XBT), a novel task for cross-modal retrieval that focuses on the compatibility between image and text embeddings of different Vision-Language Pretraining (VLP) models. We proposed an efficient solution using a text-only pretrained projection module, $\phi$, to align the new model’s embeddings with those of the old model, thereby enhancing training efficiency. By integrating parameter-efficient training schemes into the XBT framework, we were able to accelerate the model’s training while preserving the original VLP’s zero-shot capabilities. Our approach, demonstrated on various cross-modal benchmarks, effectively builds cross-modal retrieval systems without backfilling, offering an efficient and environmentally friendly solution in response to the VLP improvements.
Appendix A Appendix
A.1 Dataset Examples, More Visualization
@@
A.2 More Implementation Details
Training Recipe
A.3 Further Analysis & Discussion
@@
Continual learning.
In line with the literature on BT works [33, 28], we set up a continual learning scenario for a sequence of model updates (old model B32, new model L14, and better new model H14) and present the retrieval results in Table 6. Initially, we apply XBT to L14 to ensure compatibility with B32. Subsequently, we apply XBT to H14 to ensure compatibility with the previously learned L32, which is already compatible with B32. For this process, we divide both the image-text pairs and the text-only pretraining train sets into two halves, using each separately for each case. A comparison with the original results reported in Table 6 confirms that our XBT performs well in the continual learning scenario either. It achieves cross-modal backward compatibility while leveraging the power of the improved new model.
@@
Research question: image-only pretraining.
There exists a possibility of training $\phi$ using only images, which would replace the text-only pretraining stage. We opted for the text-only scheme because text data is easier to obtain than image data and can provide explicit semantics. These include abstract concepts, logical reasoning, and specific details that may not be visually represented. Text can express nuances of language, sentiment, and context that can be challenging to depict in an image.
However, images can also convey visual semantics that are difficult to describe in words. They can provide a wealth of details about the visual world, including color, shape, texture, spatial relationships, and visual patterns. Images can also convey non-verbal information, such as emotions and actions, that can be difficult to express in text. Therefore, pretraining $\phi$ only with images and comparing it with the text-only approach would be an interesting direction for future research.

### v2

@@ 1 Introduction
With these generated synthetic old image and text embeddings, which can be considered as aligned with the new embeddings, we aim to fine-tune the new model. In this process, far fewer supervised image-text pairs are required than in the original training dataset of the VLP. This also allows us to avoid using old model parameters during training, thereby enhancing training efficiency. In addition, we incorporate parameter-efficient fine-tuning schemes [20, 19, 12] into the new model, which add small trainable parameters and do not harm any of the new model’s original parameters during training. This not only accelerates the new model’s fine-tuning process but also allows for an easy reversion to the original new-to-new retrieval by simply removing the additional parameters.
@@ 2 Related Works
The concept of Backward-compatible Training (BT) was first introduced in the study ([34]). This approach influences a new model with the learned classifier of the old model, thereby achieving backward compatibility between the old and new models. However, BT can degrade the original representational performance of the new model. To overcome this, [26] proposed aligning class-wise centers presented by the old and new models. Another approach to achieve backward compatibility is to use an additional projection to map the old embedding into the new embedding space by adding a lightweight transformation, as suggested in [13, 38]. The work of [29] further adds an auxiliary feature in preparation for future updates, while [47] uses additional dimensionality in the embedding to obtain compatibility. An online strategy that backfills the gallery on the fly is explored in [43], and [42] addresses the model regression problem. Despite this progress, the scenario considering cross-modal retrieval between image and text, which has many real-world applications, remains unexplored. Our study on XBT in this paper aims to fill this gap.
Vision-Language Continual and Transfer Learning
@@
The fields of continual learning [2, 5, 7] and transfer learning [25, 48] bear similarities to backward-compatible representation learning, as all aim to update an existing model to boost performance. In the realm of multi-modal representation learning for cross-modal retrieval, continual learning approaches like [40] propose methods to prevent catastrophic forgetting across different modalities. Transfer learning approaches like [45] suggest strategies for transferring knowledge from previously labeled categories (source modality) to new, unlabeled categories (target modality). However, our proposed XBT stands apart from these methods as it specifically tackles the challenge of maintaining backward compatibility between old and new models. This unique attribute makes XBT ideally suited for backfill-free retrieval scenarios, where the embeddings of the old model remain unchanged, yet we can still leverage the enhanced performance of the new model.
3 Method
@@
Our goal is to construct a backfill-free, embedding-based, cross-modal retrieval system using VLP models, which are configured with an image encoder $E^{I}$ and a text encoder $E^{T}$. When a new better performing VLP model $\{E^{I}_{new},E^{T}_{new}\}$ emerges, we aim to ensure its compatibility with the old model $\{E^{I}_{old},E^{T}_{old}\}$ that was used to construct the gallery. To achieve this, we introduce Cross(X)-modal Backward-compatible Training (XBT).
Given an image $x$ and text caption $t$, XBT enables retrieval between a new image embedding $v_{new}=E^{I}_{new}(x)$ and the text embeddings ($w_{old}$), as well as between a new text embedding $w_{new}=E^{T}_{new}(t)$ and the image embeddings ($v_{old}$). We denote backward compatible embeddings as $\bar{v}$ and $\bar{w}$. All embeddings we utilize in this work are l2-normalized.
@@ 3.1 Criterion for Cross-modal Backward Compatibility
$d(w_{new_{i}},v_{old_{j}})\leq d(w_{old_{i}},v_{old_{j}}),\forall y_{i}=y_{j},$ (1)
$d(w_{new_{i}},v_{old_{j}})\geq d(w_{old_{i}},v_{old_{j}}),\forall y_{i}\neq y_{j},$
$d(v_{new_{i}},w_{old_{j}})\leq d(v_{old_{i}},w_{old_{j}}),\forall y_{i}=y_{j},$
$d(v_{new_{i}},w_{old_{j}})\geq d(v_{old_{i}},w_{old_{j}}),\forall y_{i}\neq y_{j},$
@@ 3.2 Text-only Pretraining
Figure 2: An illustration of the text-only pretraining of $\phi$ (above), and XBT with $\phi$ (below). Using only text samples, $\phi$ is trained to approximate distribution of old text embeddings from that of new ones. After pretraining, the same $\phi$ is used to generate both of synthetic old image and text embeddings from the new VLP embeddings and train to learn cross-modal backward-compatible representation.
@@ 4.1 Setup
Model Training.
Our XBT process involves two distinct training stages: text-only pretraining (See Sec. 3.2) and image-text supervised training (See Sec. 3.3). For the text-only pretraining stage, we utilize the text samples from a subset of the 115M filtered web-collected image-text paired dataset from BLIP [21], comprising around 67M available pairs (58.2% of the total). We construct a text corpus $D_{T}$ in Eqn. 3 using these text samples. For the subsequent image-text supervised training stage, we use a smaller subset of 4M image-text pairs from the same dataset to construct a supervised dataset $D$ in Eqn. 5. It is important to note that these subsets are significantly smaller than the dataset scale (400M, 2B or more) used to build CLIP models [28, 32].
Model Evaluation.
We validate the effectiveness of XBT with a benchmark comprising three popular image-text paired datasets for cross-modal retrieval evaluation. The first is the nocaps dataset [1], which offers a diverse distribution of object names and facilitates a more detailed analysis. We employ the validation split of this dataset, which consists of images each paired with 10 relevant textual captions, totaling 4,500 images and 45,000 captions.
To evaluate XBT at a larger scale, we also include the Flickr [14] and COCO [23] datasets, which consist of images that are each paired with 5 relevant textual captions. For the Flickr dataset, we use the entire dataset, encompassing 31,783 images and 158,915 captions. For the COCO dataset, we use the validation split, which includes 35,136 images and 175,680 captions. We notate the two datasets as Large-Flickr and Large-COCO respectively for the remainder of the paper.
@@ 4.2 Main Results
Table 2: Cross-modal retrieval results on Large-Flickr and Large-COCO.
Dataset Old Model / New Model Method Text $\text{Text Query}(w,\bar{w})/{\text{Image Gallery}}(v)$ Image $\text{Image Query}(v,\bar{v})/{\text{Text Gallery}}(w)$ Case R@1 R@5 R@10 R@50 Case R@1 R@5 R@10 R@50 Large -Flickr Original CLIP-ViT-B32 - $w/v$ 21.54 41.24 50.72 72.69 $v/w$ 40.35 64.42 73.39 89.91 Cross-modal Backward Compatible Training CLIP-ViT-B32 / CLIP-ViT-L14 Base 19.02 37.25 46.63 69.32 36.74 60.61 70.22 88.34 XBT $\bar{w}_{new}/v_{old}$ 22.38 42.71 52.41 74.80 $\bar{v}_{new}/w_{old}$ 42.47 66.04 74.87 90.79 Base 32.29 55.18 64.68 83.60 36.10 59.80 69.95 89.01 XBT $\bar{w}_{new}/\bar{v}_{new}$ 39.59 62.89 71.75 88.13 $\bar{v}_{new}/\bar{w}_{new}$ 43.50 68.50 77.99 93.21 CLIP-ViT-B32 / CLIP-ViT-H14 Base 20.27 39.17 48.59 70.67 35.53 59.44 69.06 87.67 XBT $\bar{w}_{new}/v_{old}$ 23.70 44.47 54.19 76.03 $\bar{v}_{new}/w_{old}$ 41.09 65.18 74.32 90.16 Base 32.94 55.94 65.26 83.77 34.81 59.55 69.96 88.81 XBT $\bar{w}_{new}/\bar{v}_{new}$ 40.42 63.90 72.74 88.47 $\bar{v}_{new}/\bar{w}_{new}$ 46.35 71.18 80.01 94.08 Large -COCO Original CLIP-ViT-B32 - $w/v$ 14.44 30.20 38.96 62.30 $v/w$ 28.62 50.17 59.67 80.22 Cross-modal Backward Compatible Training CLIP-ViT-B32 / CLIP-ViT-L14 Base 12.97 28.14 36.48 59.78 26.32 46.91 56.54 77.86 XBT $\bar{w}_{new}/v_{old}$ 15.55 32.27 41.34 65.11 $\bar{v}_{new}/w_{old}$ 30.73 52.58 62.30 81.83 Base 22.16 41.65 50.98 73.06 27.26 48.16 57.80 79.57 XBT $\bar{w}_{new}/\bar{v}_{new}$ 27.61 48.69 58.08 78.83 $\bar{v}_{new}/\bar{w}_{new}$ 32.92 55.37 64.89 84.67 CLIP-ViT-B32 / CLIP-ViT-H14 Base 13.88 29.55 38.23 61.57 26.19 46.75 56.30 77.96 XBT $\bar{w}_{new}/v_{old}$ 15.82 32.79 41.84 65.45 $\bar{v}_{new}/w_{old}$ 30.33 51.28 59.97 81.66 Base 23.20 43.22 52.70 74.68 27.31 48.56 58.60 80.09 XBT $\bar{w}_{new}/\bar{v}_{new}$ 28.14 49.48 58.88 79.32 $\bar{v}_{new}/\bar{w}_{new}$ 34.54 57.29 66.79 85.92
@@
[…]
Large Scale Retrieval. In this experiment, we assess the performance of cross-modal backward compatibility in more practical use cases. We use larger datasets, namely Large-Flickr and Large-COCO, as shown in Table 2. These datasets provide more than six times the number of samples typically used in image-text retrieval literature [37, 6], which make the retrieval process more challenging. We use the second-best performing Base method from Table 1 as a basis for comparison. To validate the applicability of XBT across various VLP models, we utilize three CLIP baselines: B32, L14, and H14. Additionally, we explore different combinations, such as B32/L14 and B32/H14, to further analyze the effectiveness of XBT, beyond the comparisons made in Table 1 (B32/L14 and L14/H32).
@@ 4.3 Further Analysis
Continual learning. In line with the literature on BT works [34, 29], we set up a continual learning scenario for a sequence of model updates (old model B32, new model L14, and better new model H14) and present the retrieval results in Table 3. Initially, we apply XBT to L14 to ensure compatibility with B32. Subsequently, we apply XBT to H14 to ensure compatibility with the previously learned L32, which is already compatible with B32. For this process, we divide both the image-text pairs and the text-only pretraining train sets into two halves, using each separately for each case. A comparison with the original results reported in Table 3 confirms that our XBT performs well in the continual learning scenario either. It achieves cross-modal backward compatibility while leveraging the power of the improved new model.
Ablation Study. To validate XBT, we conduct further analysis as shown in Table 4. By comparing (a) and (b), we observe that introducing noise during $\phi$ training aids in generalization. The comparison between (c) and (d) demonstrates that the scale of $D_{T}$ is important, supporting our assumption that a sufficient number of text samples can help build a robust $\phi$. (e) outperforms (a), confirming that utilizing more image-text pairs can enhance XBT. The lower performance of (f) also aligns with the notion that the number of image-text pairs is crucial. In (g), when we replace $D$ with CC3M [33], from which we can obtain around 2.4M, the performance is similar to (f), suggesting that XBT can be effectively applied with other datasets. In (h), we replace text-only pretraining with image-only pretraining of the same scale and observe that text-only pretraining performs better in both retrieval scenarios, demonstrating its efficiency and superiority.
5 Conclusion
In this paper, we introduced Cross-modal Backward-compatible Training (XBT), a novel task for cross-modal retrieval that focuses on the compatibility between image and text embeddings of different Vision-Language Pretraining (VLP) models. We proposed an efficient solution using a text-only pretrained projection module, $\phi$, to align the new model’s embeddings with those of the old model, thereby enhancing training efficiency. By integrating parameter-efficient training schemes into the XBT framework, we were able to accelerate the model’s training while preserving the original VLP’s zero-shot capabilities. Our approach, demonstrated on various cross-modal benchmarks, effectively builds cross-modal retrieval systems without backfilling, offering an efficient and environmentally friendly solution in response to the VLP improvements.
Supplementary Material
Different VLP models.
To further explore the capacity of the VLP model architecture’s generalization of XBT, we evaluate it using BLIP [21] Base and Large models. We employ checkpoints of Salesforce/blip-itm-base-coco and Salesforce/blip-itm-large-coco from the Huggingface library. We apply XBT on old and new BLIP models in the same fashion with our CLIP applications, and the results appear in Table 5. From the results, we confirm that XBT provides cross-modal backward-compatibility to the BLIP models too.
@@ 5 Conclusion
5.1 Dataset Examples, More Visualization
@@
5.2 Further Analysis & Discussion